OpenStudy (anonymous):

A long time ago, in a galaxy far, far away, Lama Su takes Obi-Wan Kenobi on a tour of the clone factory in Tipoca City on Kamino. In order to obtain the biomass necessary to construct the clones, a gigantic vat is filled with 700 gallons of a solution of liquefied Hutt fat containing 80 pounds of granulated Wookie hair. Then, pure liquefied Hutt fat is pumped into the vat at a rate of 6 gallons per minute, while the well-stirred mixture is removed from the vat at a rate of 9 gallons per minute. After 3 hours of pumping pure Hutt fat into the vat, how many pounds of granulated Wookie hair are left in the vat?

OpenStudy (anonymous):

who ever came up with this question should be killed,they screwed up star wars worse than Disney did making that musical

ganeshie8 (ganeshie8):

plugin the initial value, solve the constant c

OpenStudy (anonymous):

Ok here's how I did it: t=180mins Let H(t)=amount of liquefied Huttfat (in gallons) after t minutes Let W(t)=amount of wookie hair (in pounds) after t minutes \[\frac{ dw }{ dt }\]=rate in(of hair) - rate out(of hair) =[(rate in(of soln)*concentration of soln. coming in)-(rate out (of soln)*concentration of soln coming out)] =(6gal/min*0bl/gal)-(9gal/min*(hair at time t/soln of time t) =0lb/min-9gal*(W lb/ H gal) dW/dt=-9W/(700-3t)=(-9/(700-3t))*dt=dW/W \[\int\limits \frac{ dW }{W }=\int\limits \frac{ -9 }{ 700-3t }dt \rightarrow \ln \left| W \right|=-9\ln \left| 700-3t \right|+c\] \[e^{\ln \left| W \right|}=e^{-9\ln \left| 700-3t \right|} +e^{C}\] are these the right steps to find C and find the amount of wookie hair? If not what steps do I need to add or revise?

OpenStudy (anonymous):

this is the worst star wars question ever

OpenStudy (anonymous):

scary lol

ganeshie8 (ganeshie8):

it looks okay, except for a minor mistake when evaluating integral

ganeshie8 (ganeshie8):

you should get : \(\large \int\limits \frac{ dW }{W }=\int\limits \frac{ -9 }{ 700-3t }dt \rightarrow \ln \left| W \right|=\dfrac{-9\ln \left| 700-3t \right|}{-3}+c\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

Oh yah and the 9 cancelles out 3 yah I see

ganeshie8 (ganeshie8):

yes and it simplifies to the wolfram solution click my first reply